OpenStudy (anonymous):

use laws of log to combine. log12+1/2log5-log2

OpenStudy (anonymous):

Here are the laws, and you can apply them: log(a)+log(b) = log(a*b) log(a)-log(b) = log(a/b) a*log(b) = log(b^a)

OpenStudy (anonymous):

It decomposes to log(12 * 5^(1/2) / 2).

OpenStudy (anonymous):

Which breaks down further to be log(6*5^(1/2)).

OpenStudy (anonymous):

\[\log(6 \sqrt(5))\]

OpenStudy (anonymous):

okay thanks, I did decompose it to log 12*5^1/2)/2 but i didn't break it down, i understand my fault